[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Oct 09, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Oct 9 16:22:04 UTC 2017


Here is today's summary of NHC computer operations:
--- PS

----------

TAFB:
1.) Atlantic Graphicast is still fuzzy when sent from Ballyhoo. Retrieved  
file on Grunt and sent it from there. Atlantic graphicast now appears clear  
on the NHC web page - Matt Sardi already working with Doug on this. This is  
likely related to NAWIPS/REHL7 issue. A decision based on what NCO can or  
not do about this with NAWIPS will be made this week as to whether we  
revert back to REL6.

2.) Scatterometry is not currently displaying at any TAFB workstation.  
Imagery appeared to stop displaying around 0415Z on 7 Oct. Am referring to  
Manati site for the latet imagery. This appears to have been a NOAA wide  
issue with ASCAT per SDM. TSB will check on this on Tuesday.
HSU:
1.) New SSU NAWIPS workstation (fishbowl) suddenly stopped working and  
logged itself out. SSU reports this is the 2nd time it has happened. They  
are not using it for operational products for this reason since it doesn't  
seem to be stable - Doug Gaer will investigate this on Tuesday. NCO ticket  
will be open accordingly if needed.

2.) We need a full time person just to maintain National TCV. Folks this  
will be addressed. Yes it is clear we need someone to coordinate on this  
end not just during day and evening but also overnight. This is a  
coordination challenge not a software one. But given the challenges this  
season it has worked rather well revealing along the way some nice side  
benefits. This is a discussion not for this log but for the off season.  
Proper pre TCV procedures will also need to be part of this discussion ;)


3.) Bonefish froze and X-server had to be restarted. Satellite loops  
continued to load slowly after that. Doug Gaer will investigate on Tueday.


4.) Not much of an issue, but just wanted to pass on that the aids colors  
are not the standard ones. TSB/Doug Gaer to investigate in case it is  
related to the upgrade and new OS.
Other items:
1. HFIP Writing Team Meeting Wed and Thursday.
NHC RFCs:
1.) RFC 3812 - NHC Memory Upgrades on ATFC Server 1 scheduled for Thursday  
10/12.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171009/8bcf70cc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list